Status Report

NASA ISS On-Orbit Status 26 July 2008

By SpaceRef Editor
July 26, 2008
Filed under , , ,
NASA ISS On-Orbit Status 26 July 2008

All ISS systems continue to function nominally, except those noted previously or below. Saturday – mostly off duty for CDR Volkov, FE-1 Kononenko & FE-2 Chamitoff.

Sergey Volkov began his workday by attending to the current experiment session with the Russian/German TEKh-20 Plasma Crystal-3+ (Plazmennyi-Kristall/PK-3+) payload, activating the turbopump in the Service Module (SM)’s Transfer Compartment (PkhO) for keeping the vacuum chamber (ZB) in the SM Work Compartment (RO) evacuated. The turbopump will be deactivated again tonight at ~5:25pm EDT before sleeptime. [Main objective of PK-3 is to study dust plasma wave propagation and dispersion ratio at a specified power of HF discharge, pressure, and a varied number of particles.]

The crew performed the regular weekly three-hour task of thorough station cleaning. ["Uborka", usually done on Saturdays, includes removal of food waste products, cleaning of compartments with vacuum cleaner, damp cleaning of the SM dining table, other frequently touched surfaces and surfaces where trash is collected, as well as the FE’s sleep station with a standard cleaning solution; also, fan screens and grilles are cleaned to avoid temperature rises. Special cleaning is also done every 90 days on the HEPA (high-efficiency particulate air) bacteria filters in the Lab.]

As part of the house cleaning, the crew also conducted regular maintenance inspection & cleaning on fan grilles in the FGB (TsV2), SM (VPkhO, VPrK, FS5, FS6 & FS9), DC1 (V3) screens and PF1/PF2 dust filters, as well as in the COL (Columbus Orbital Laboratory).

Gregory filled out the regular FFQ (Food Frequency Questionnaire), his sixth, on the MEC (Medical Equipment Computer). [On the FFQs, NASA astronauts keep a personalized log of their nutritional intake over time on special MEC software. Recorded are the amounts consumed during the past week of such food items as beverages, cereals, grains, eggs, breads, snacks, sweets, fruit, beans, soup, vegetables, dairy, fish, meat, chicken, sauces & spreads, and vitamins. The FFQ is performed once a week to estimate nutrient intake from the previous week and to give recommendations to ground specialists that help maintain optimal crew health. Weekly estimation has been verified to be reliable enough that nutrients do not need to be tracked daily.]

At ~9:45am EDT, the crewmembers conducted their regular WPC (Weekly Planning Conference) with the ground, discussing next week’s "Look-Ahead Plan" (prepared jointly by MCC-H and TsUP timeline planners) via S-band/audio, reviewing the monthly calendar, upcoming activities, and any concerns about future on-orbit events.

Conducting his VolSci voluntary weekend science program, the FE-2 set up for several hours of a SHERE (Shear History Extension Rheology Experiment) session with the CGBA (Commercial Generic Bioprocessing Apparatus) activating the MSG (Microgravity Science Glovebox) with its A31p laptop, powered on the SHERE payload in the CGBA, unstowing Fluid Module #30 and performed a dry run, followed by CGBA power down, data transfer and deactivation of SHERE and the MSG via the A31p.

Afterwards, Gregory took measurements for the regular atmospheric status check for ppCO2 (Carbon Dioxide partial pressure) in the Lab, SM (at panel 449) and COL (Columbus Orbital Laboratory), using the hand-held CDMK (CO2 Monitoring Kit, #1002). [The CDMK was then deactivated and returned to its stowage location at LA B1S2.]

Oleg completed the routine daily servicing of the SOZh system (Environment Control & Life Support System, ECLSS) in the Service Module (SM). [Regular daily SOZh maintenance consists, among else, of checking the ASU toilet facilities, replacement of the KTO & KBO solid waste containers and replacement of EDV-SV waste water and EDV-U urine containers.]

The three crewmembers conducted their regular 2.5-hr. physical workout program (about half of which is used for setup & post-exercise personal hygiene) on the CEVIS cycle ergometer (FE-2), TVIS treadmill (CDR, FE-1), RED resistive exercise device (CDR, FE-1, FE-2), and VELO bike with bungee cord load trainer (CDR, FE-1). Later, Greg transferred the exercise data file to the MEC (Medical Equipment Computer) laptop for downlink, including the daily wristband HRM (Heart Rate Monitor) data of the workouts on RED, followed by their erasure
on the HRM storage medium (done six times a week).

The two cosmonauts had their weekly PFCs (Private Family Conferences) via S-band/audio and Ku-band/MS-NetMeeting application (which displays the uplinked ground video on an SSC laptop), Sergey at ~8:20am, Oleg at ~10:20am.

Weekly Science Update (Expedition Seventeen — Week 14)

3-D SPACE: Third session for Greg Chamitoff is currently planned for 7/30.

ALTCRISS (Alteino Long Term monitoring of Cosmic Rays on the ISS): Measurements continue in the FGB module. Last memory card exchange was performed on 7/1. Next activities are planned on 7/22 for the rotation of the ALTEINO instrument and the exchange of the memory card. Subsequent data downlink is currently planned on 7/23.

ANITA (Analyzing Interferometer for Ambient Air): Continuing.

BCAT-3 (Binary Colloidal Alloy Test 3): Reserve.

CARDIOCOG-2: Completed.

CCISS (Cardiovascular & Cerebrovascular Control on Return from ISS): Reserve.

CFE (Capillary Flow Experiment): Reserve.

CW/CR (Cell Wall/Resist Wall) in EMCS (European Modular Cultivation System): Samples returned on 1J.

CSI-2/CGBA (CGBA Science Insert #2/Commercial Generic Bioprocessing Apparatus): The Silicate Garden Hab S/N2 ground unit was activated at approximately 5:15pm EDT on 7/11. The flight unit was activated at approximately 1:25pm on 7/14 via remote command. There are 2 cameras in the flight unit imaging 2 chambers each for a total of 4 chambers on Side 1. Fibers continue to grow from the crystals in strange formations in the flight unit.

CGBA-2 (Commercial Generic Bioprocessing Apparatus 2): Complete.

CSLM-2 (Coarsening in Solid-Liquid Mixtures 2): Complete.

EarthKAM (Earth Knowledge Acquired by Middle School Students): Complete.

ELITE-S2 (Elaboratore Immagini Televisive – Space 2): Planned.

EPO (Educational Payload Operations): Reserve.

ETD (Eye Tracking Device): Completed.

EuTEF (European Technology Exposure Facility): On 7/16, EuTEF encountered a MIL-Bus error (this happened similarly in the past). The platform has to be power-cycled and the entire EuTEF facility re-initiated. On 7/20, a similar DHPU (Data Handling Processing Unit) problem was encountered again. Although this problem does not prevent EuTEF to function, ground teams are working hard to characterize it and find a solution. A DHPU software patch is currently being developed to fix the link error issues encountered with DEBIE-2 and FIPEX instruments. — DEBIE-2: Instrument could not be restarted immediately after the DHPU problem on 7/20, and was finally re-activated on 7/22) for a short diagnostic run. EBIE-2 was stopped on 7/23 to allow science acquisition with FIPEX. Generic status: link error still in work. However, a work-around allows for regular science data acquisition using an on-board Instrument Operations Procedure (IOP), but not yet in conjunction with any FIPEX IOP – a final software patch, which will enhance the present instrument capabilities, is under finalization on ground;– DOSTEL: On-going science acquisition;– EuTEMP: Currently inactive as planned;– EVC: Inactive this week;– EXPOSE: On-going science acquisition;– FIPEX: Instrument re-activated after DHPU problem on 7/20. Science acquisition since 7/23;– MEDET: Instrument is not active due to low temperature; — PLEGPAY: Currently in READY mode, no science data acquisition on-going;– TRIBOLAB: The instrument is currently in Stand-By Mode. The Ball-Bearing experiment would be restarted towards mid-Aug. Analysis is on-going on ground.

FSL (Fluid Science Laboratory): FSL is back and running. Further FSL commissioning activities remain to be done (see following item).

GEOFLOW: On 7/23, Greg swapped the FSL optical targets, and the Optical check-out 2 has been successfully performed from ground before the ATV reboost. At this point, everything looks ready to remove this optical target and to insert the GEOFLOW Experiment Container on 7/28. After some initial tests, the GEOFLOW science program is planned to start on 7/29.

IMMUNO (Neuroendocrine & Immune Responses in Humans During & After Long Term Stay at ISS): Complete.

InSPACE-2 (Investigating the Structure of Paramagnetic Aggregates from Colloidal Emulsions 2): In progress.

Integrated Immune: In progress.

KUBIK-FM1/ KUBIK-FM2 Centrifuge/Incubators: Completed.

LOCAD-PTS (Lab-on-a-Chip Application Development-Portable Test System): Planned.

MISSE (Materials ISS Experiment): Ongoing.

MTR-2 (Russian radiation measurements): Passive dosimeters measurements in DC1 “Pirs”.

MULTIGEN-1: Completed.

MSG-SAME (Microgravity Science Glovebox): Complete.

NOA-2 (Nitric Oxide Analyzer, ESA): In progress.


PADLES (Passive Dosimeter for Lifescience Experiment in Space): In progress.

PMDIS (Perceptual Motor Deficits in Space): Complete.

SAMS/MAMS (Space & Microgravity Acceleration Measurement Systems): Ongoing.

SAMPLE: Complete.

SHERE (Shear History Extensional Rheology Experiment): The SHERE team thanks Greg for selecting SHERE for Voluntary Science checkout activities performed on GMT 202. The checkout tests were a great success and showed that the SHERE hardware is functioning properly. The SHERE team looks forward to working with Greg again tomorrow, GMT 208, to perform the SHERE Dry Run. The recent fit problem with the Fluid Module Stowage Tray is being worked and will not affect the Dry Run.

SLEEP (Sleep-Wake Actigraphy & Light Exposure during Spaceflight): The HRP team thanks Greg for completing his first week of Sleep logs and his Actiwatch download initialization session. HRP is currently targeting the week of 8/4 for Greg’s next week of Sleep logging and the week of 8/11 for his next Actiwatch download. Additional Sleep logging on non-targeted weeks is above and beyond and greatly appreciated by the PI.

SOLAR (Solar Monitoring Observatory): After a successful troubleshooting activity by Greg on 7/21 (GMT203) and subsequent downlink of the log file generated during the boot sequence of SOLAR, ground teams have been very busy analyzing this new input with some very good news: the SOLAR LAN interface has been confirmed to be fully nominal. Next troubleshooting step was to focus on the Columbus LAN Switch#2 (CLSW#2), and on 7/23 (GMT205), ground activities were performed. SOLAR power feeder #1 was re-activated immediately and was forced to communicate via the other Columbus LAN switch (CLSW#1). Immediately, SOLAR started to re-send telemetry to the ground. Full functionality of the platform has been recovered. The instruments have been progressively warming up, thanks to continuous power via feeder #1 after this harsh beta angle period. The next Sun observation window starts on 7/31.– SOVIM: Instrument confirmed fully functional; awaiting the Sun; — SOLSPEC: Instrument confirmed fully functional; awaiting the Sun;– SOLACES: Instrument confirmed fully functional; awaiting the Sun.

SOLO: Planned.

SPHERES (Synchronized Position Hold, Engage, Reorient, Experimental Satellite): In progress.

Swab (Characterization of Microorganisms & Allergens in Spacecraft): Complete.

TRAC (Test of Reaction & Adaptation Capabilities): Planned.


WAICO #1 (Waving and Coiling of Arabidopsis Roots at Different g-levels): The four ECs (Experiment Containers) of Rotor B are yet to be retrieved by the crew from the blocked Rotor B. BIOLAB could not be activated from ground yet, pending resolution of the Smoke Detector issue. A troubleshooting plan has been developed to recover all the BIOLAB functionalities. The first step is to exchange an ESEM Power Board, in which this activity is planned later on today, i.e. 7/25. Hopefully, this will resolve the Smoke Detector issue and the ground will proceed with the next steps; starting with the BIOLAB activation from ground on 7/26. The remaining WAICO-#1 ECs will be disposed and replaced by 6 Reference ECs on Rotor B. The remaining WAICO-#1 ECs will be disposed and replaced by six Reference ECs on Rotor B.

CEO (Crew Earth Observations): Through 7/15 the ground has received a total of 3,391 frames of CEO images for review and cataloging. New imagery received this week with times corresponding to CEO target request times included: Lake Poopo, Bolivia (interesting new imagery of the target is now under review); Jarvis Island, equatorial Pacific (only a solitary, late-shot frame of this isolated island); and Moorea Coral Reef, Tahiti (good imagery acquired, a bit late-oblique, and a touch soft). “As you are aware CEO target opportunities have been limited by low light and winter weather. It will take good fortune to spot even a large, bright iceberg under those conditions. We hope you may have better luck with the Polar Mesospheric Cloud opportunities. This lull is allowing us to work through our backlog. You shot excellent imagery of: Mount Rainer in with snow, the Red River Valley, the Tigris-Euphrates Delta, and much of the Afar Rift Zone target. Good work! We suggest technique-wise that you continue to practice with the long lens settings to improve consistency of focus, and continue to practice providing an overlap between successive images when mapping a feature or a regional target. The composition and quality of your images is high and we do appreciate you all your efforts.”

CEO photography can be studied at this “Gateway” website: (as of 3/1/08, this database contained 757,605 views of the Earth from space, with 314,000 from the ISS alone).

Significant Events Ahead (all dates Eastern Time, some changes possible.):
08/30/08 — Progress M-64/29P undocking, from FGB nadir
09/05/08 — ATV1 undocking, from SM aft port (loiter until ~9/25 for nighttime reentry/observation)
09/10/08 — Progress M-65/30P launch
09/12/08 — Progress M-65/30P docking (SM aft port)
10/01/08 — NASA 50 Years
10/08/08 — STS-125/Atlantis Hubble Space Telescope Service Mission 4 (SM4)
10/11/08 — Progress M-65/30P undocking (from SM aft port)
10/12/08 — Soyuz TMA-13/17S launch
10/14/08 — Soyuz TMA-13/17S docking (FGB nadir port)
10/23/08 — Soyuz TMA-12/16S undocking (DC1 nadir)
11/10/08 — STS-126/Endeavour/ULF2 launch – MPLM Leonardo, LMC
11/12/08 — STS-126/Endeavour/ULF2 docking
11/20/08 — ISS 10 Years
11/26/08 — Progress M-66/31P launch
11/28/08 — Progress M-66/31P docking
02/10/09 — Progress M-67/32P launch
02/12/09 — Progress M-67/32P docking
02/12/09 — STS-119/Discovery/15A launch – S6 truss segment
03/25/09 — Soyuz TMA-14/18S launch
05/15/09 — STS-127/Endeavour/2J/A launch – JEM EF, ELM-ES, ICC-VLD
07/30/09 — STS-128/Atlantis/17A – MPLM(P), last crew rotation
05/27/09 — Six-person crew on ISS (following Soyuz 19S docking, May ’09)
10/15/09 — STS-129/Discovery/ULF3 – ELC1, ELC2
12/10/09 — STS-130/Endeavour/20A – Node-3 + Cupola
02/11/10 — STS-131/Atlantis/19A – MPLM(P)
04/08/10 — STS-132/Discovery/ ULF4 – ICC-VLD, MRM1
05/31/10 — STS-133/Endeavour/ULF5 – ELC3, ELC4 (contingency).

SpaceRef staff editor.